determine whether the pair of fractions are proportional.
7/9 = 21/27?

evaluate the expression then simplify. this is a fraction .
12divided by(3).(-8)
(3 to 2ndpower + 1 to 2nd power)+2
: determine whether the pair of fractions are proportional.
7/9 = 21/27?

evaluate the expression then simplify. this is a fraction .
12divided by(3).(-8)
(3 to 2ndpower + 1 to 2nd power)+2

Answer by Fombitz(1755)

To determine proportionality, cross multiply.
If the products are equal, the fractions are proportional.
7/9 = 21/27
7(27)=9(21)
189=189
Yes, they are proportional.
(7/9)(3/3)=21/27
.
.
.
12/(3*(-8))=12/(-24)=-(1/2)
.
.
.
Remember the order of operations: parentheses, powers, then addition.
(3^2+1^2)+2=(9+1)+2
(3^2+1^2)+2=(10)+2
(3^2+1^2)+2=12